The area of a parallelogram is 45 sq units. Base is 2x+1. Height x+1. What is the base & height in units?
Answer by Boreal(15235)

Area=bh=2x^2+3x+1, multiplying the two
that equals 45
2x^2+3x-44=0
(2x+11)(x-4)=0
x=4, only positive root
base is 9, height is 5
product is 45.